A bike successfully made a turn around a circular path whose radius is 13.0 m. At what angle is the road banked if its velocity was 12.0 m/s/ 




1
Expert's answer
2021-12-02T10:07:23-0500

"r = 13m"

"v = 12\frac{m}{s}"

"m -\text{bike weight}"

"\alpha - \text{ banked angle}"

"\text{Forces acting on a bike:}"

"F_c-\text{the centripetal force}"

"F_c = ma"

"a =\frac{v^2}{r}"

"F_c= m\frac{v^2}{r}"

"N - \text{road reaction force}"

"F_g -\text{gravity}"

"F_g = mg"

"\text{Friction force is not specified in the problem statement}"

"\text{Vertical force projection}"

"N\cos\alpha-F_g=0"

"N\cos\alpha= mg(1)"

"\text{Horisontal force projection}"

"F_c-N\sin\alpha=0"

"N\sin\alpha=m\frac{v^2}{r}(2)"

"\text{From }(1) \text{ and }(2)"

"\tan \alpha= \frac{v^2}{rg}= \frac{12^2}{13*9.8}\approx1.1303"

"\arctan(1.1303)=48.5\degree"


"\text{Answer: } 48.5\degree"